# Decentralized Optimization with Coupled Constraints

### v3

@@ 1 Introduction
Coupled constraints arise in various application scenarios, where sharing of resources or information takes place. Often, due to the distributed nature of such problems, decentralization is desired for communication and/or privacy related reasons. Let us briefly describe several practical cases of optimization problems with coupled constraints.
@@
To handle the consensus constraint, decentralized algorithms either reformulate it as $\sum_{i=1}^{n}\mathbf{W}_{i}x_{i}=0$, where $\mathbf{W}_{i}$ is the $i$-th vertical block of a gossip matrix (an example of which is the communication graph’s Laplacian), or utilize the closely related mixing matrix approach Gorbunov et al. (2022). Mixing and gossip matrices are used because they are communication-friendly: calculating the sum $\sum_{i=1}^{n}\mathbf{W}_{i}x_{i}$ only requires each compute node to communicate once with each of its adjacent nodes. Clearly, consensus optimization with gossip matrix reformulation can be reduced to (1) by setting $\mathbf{A}_{i}=\mathbf{W}_{i}$. However, the principial difference between this example and (1), is that (1) does not assume $\mathbf{A}_{i}$ to be communication-friendly.
@@ 2 Related work and our contribution
Decentralized optimization algorithms were initially proposed for consensus optimization Nedić and Ozdaglar (2009), based on earlier research in distributed optimization Tsitsiklis (1984); Bertsekas and Tsitsiklis (1989) and algorithms for decentralized averaging (consensus or gossip algorithms) Boyd et al. (2006); Olshevsky and Tsitsiklis (2009), which assumed the existence of a communication network, as does the present paper. The optimal (accelerated) $O(\sqrt{\kappa_{f}}\ln\frac{1}{\varepsilon})$ complexity for consensus optimization was first achieved with a dual (requiring to compute gradients of Fenchel conjugates of $f_{i}(x)$) accelerated gradient descent in Scaman et al. (2017), where the corresponding complexity lower bounds were also established. This result was later generalized to primal algorithms (which use gradients of the functions $f_{i}(x)$ themselves), time-varying communication graphs Li and Lin (2021); Kovalev et al. (2021) and stochastic gradients Dvinskikh and Gasnikov (2021). Today there also exist algorithms with communication compression Beznosikov et al. (2023), asynchronous algorithms Koloskova (2024), algorithms for saddle-point formulations Rogozin et al. (2021) and gradient-free oracles Beznosikov et al. (2020), making decentralized consensus optimization a quite well-developed field Nedić (2020); Gorbunov et al. (2022), benefiting systems control Ram et al. (2009) and machine learning Lian et al. (2017).
Beginning with the addition of local constraints to consensus optimization Nedic et al. (2010); Zhu and Martinez (2011), constrained decentralized optimization has been established as a research direction. A zoo of distributed problems with constraints was investigated in Necoara et al. (2011), see also Necoara and Nedelcu (2014, 2015).
Primarily motivated by the demand from the power systems community, various decentralized algorithms for coupled constraints have been proposed. Generally designed versatile for engineering applications, many of these algorithms assume restricted function domains Wang and Hu (2022); Liang et al. (2019); Nedić et al. (2018); Gong and Zhang (2023); Zhang et al. (2021); Wu et al. (2022), nonlinear inequality constraints Liang et al. (2019); Gong and Zhang (2023); Wu et al. (2022), time-varying graphs Zhang et al. (2021); Nedić et al. (2018) or utilize specific problem structure Wang and Hu (2022).
@@
Works of Doan and Olshevsky (2017); Li et al. (2018); Nedić et al. (2018) focus on the resource allocation problem. For undirected time-varying graphs Doan and Olshevsky (2017) proposes a first-order algorithm with $N=O(\frac{\kappa_{f}n^{2}}{B}\ln\frac{1}{\varepsilon})$ complexity bound, where $B$ is the time the graph takes to reach connectivity. Li et al. (2018) applies a combination of gradient tracking and push-sum approaches from Nedic et al. (2017) to obtain linear convergence on directed time-varying graphs in the restricted domain case, i.e., $x_{i}\in\Omega_{i}$, where $\Omega_{i}$ is a nonempty closed convex set. Nedić et al. (2018) achieve accelerated linear convergence via a proximal point method in the restricted domain case. When $\Omega_{i}=\mathbb{R}^{d}$, they also show that Nesterov’s accelerated gradient descent can be applied to achieve optimal $O(\sqrt{\kappa_{f}}\ln\frac{1}{\varepsilon})$ complexity. In Gong and Zhang (2023) an inexact proximal-point method is proposed to solve problems with coupled affine equality and convex inequality constraints. Linear convergence is proved when the inequalities are absent, and $\Omega_{i}$ are convex polyhedrons. The papers Wu et al. (2022), Chang (2016), Falsone et al. (2020) present algorithms with sub-linear convergence.
As summarized in Table 1, there is no accelerated linearly convergent algorithms for general affine-equality coupled constraints. Also, most of the algorithms require proximal oracle, which allows to handle more general problem formulations, but has higher computational burden than the first-order oracle. We propose a new first-order decentralized algorithm with optimal (accelerated) linear convergence rate. We prove its optimality by providing lower bounds for the number of objective’s gradient computations, matrix multiplications and decentralized communications, which match complexity bounds for our algorithm.
@@ 3 Mathematical setting and assumptions
By $\kappa_{f}$ we denote the condition number
$\kappa_{f}=\frac{L_{f}}{\mu_{f}}.$ (4)
@@
where the matrix $\mathbf{S}\in\mathbb{R}^{m\times m}$ is defined as $\mathbf{S}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{A}_{i}\mathbf{A}_{i}^{\top}$. We also define the condition number of the matrix $\mathbf{A}$ as $\kappa_{\mathbf{A}}=L_{\mathbf{A}}/\mu_{\mathbf{A}}$.
@@
We also assume the existence of a so-called gossip matrix $W\in\mathbb{R}^{n\times n}$ associated with the communication network $\mathcal{G}$, which satisfies certain assumptions.
@@ 4.1 Strongly convex communication-friendly reformulation
and multiplication of a vector $y=(y_{1},\ldots,y_{n})\in(\mathbb{R}^{m})^{n}$ by $\mathbf{W^{\prime}}$ can be performed efficiently in the decentralized manner if its $i$-th block component $y_{i}$ is stored at $i$-th node of the computation network. Similarly to eq. 7, we define
@@
Due to the definition of $\mathbf{W}$ and eq. 8, the simplest choice for $\mathbf{W^{\prime}}$ might be to set $\mathbf{W^{\prime}}=\mathbf{W}$. Later we will specify another way to choose $\mathbf{W^{\prime}}$ for optimal algorithmic performance.
@@
From the definitions of $\mathbf{A}$, $\mathbf{b}$ and $\mathcal{L}_{m}^{\bot}$ (eq. 3) it is clear that $\sum_{i=1}^{n}(\mathbf{A}_{i}x_{i}-b_{i})=0$ if and only if $\mathbf{A}x-\mathbf{b}\in\mathcal{L}_{m}^{\perp}$. Since $\operatorname{\mathrm{range}}\mathbf{W^{\prime}}=\mathcal{L}^{\perp}_{m}$, the constraint in problem (11) is equivalent to the coupled constraint in (1). For all $x,y$ satisfying the constraint, the augmented objective function $G(x,y)$ is equal to the original objective function $F(x)$. Therefore, problem (11) is equivalent to problem (1).
The following Lemma 1 shows that the function $G(x,y)$ is strongly convex and smooth.
@@
Let the matrix $\mathbf{B}\in\mathbb{R}^{mn\times(d+mn)}$ be defined as $\mathbf{B}=\left[\begin{matrix}\mathbf{A}&\gamma\mathbf{W^{\prime}}\end{matrix}\right]$. The following Lemma 2 connects the spectral properties of $\mathbf{B}$, $\mathbf{A}$ and $\mathbf{W^{\prime}}$.
@@
Proofs of Lemma 1 and Lemma 2 are provided in Appendix A.
@@ 4.2 Chebyshev acceleration
Chebyshev acceleration allows us to decouple the number of computations of the objective’s gradient $\nabla F(x)$ from the properties of the communication network and the constraint matrix — specifically, from the condition numbers $\kappa_{\mathbf{W}}$ and $\kappa_{\mathbf{A}}$.
First, we choose $\mathbf{W^{\prime}}$ as a polynomial $P_{\mathbf{W}}(\mathbf{W})$ to make its condition number $\kappa_{\mathbf{W^{\prime}}}$ be bounded by some small constant. Require $P_{\mathbf{W}}(\lambda)=0\Leftrightarrow\lambda=0$ for every eigenvalue $\lambda$ of $\mathbf{W}$. Then, by eq. 8, $\operatorname{\mathrm{range}}P_{\mathbf{W}}(\mathbf{W})=\operatorname{\mathrm{range}}\mathbf{W}=\mathcal{L}_{m}^{\bot}$, and we can set $\mathbf{W^{\prime}}=P_{\mathbf{W}}(\mathbf{W})$.
Next, we aim to reduce $\kappa_{\mathbf{B}}$. Let $P_{\mathbf{B}}$ be a polynomial such that $P_{\mathbf{B}}(\lambda)=0\Leftrightarrow\lambda=0$ for every eigenvalue $\lambda$ of $\mathbf{B}^{\top}\mathbf{B}$. For compactness, we denote by $u=(x,y)$ the stacked primal variable of problem (11). We denote any feasible point for the affine constraint $\mathbf{B}u_{0}=\mathbf{b}$ as $u_{0}$. Then,
$\mathbf{B}u=\mathbf{b}$ $\Leftrightarrow\mathbf{B}(u-u_{0})=0\overset{(a)}{\Leftrightarrow}\mathbf{B}^{\top}\mathbf{B}(u-u_{0})=0$
$\overset{(b)}{\Leftrightarrow}P_{\mathbf{B}}(\mathbf{B}^{\top}\mathbf{B})(u-u_{0})=0\overset{(c)}{\Leftrightarrow}\sqrt{P_{\mathbf{B}}(\mathbf{B}^{\top}\mathbf{B})}(u-u_{0})=0$
[…]
Following Salim et al. (2022a) and Scaman et al. (2017), we choose $P_{\mathbf{W}},P_{\mathbf{B}}$ to be the translated and scaled Chebyshev polynomials, because they are the best at compressing the spectrum Auzinger and Melenk (2011).
@@
Consider a matrix $\mathbf{M}$. Let $n=\left\lceil\sqrt{\frac{L_{\mathbf{M}}}{\mu_{\mathbf{M}}}}\right\rceil\geq\left\lceil\sqrt{\frac{\lambda_{\max}(\mathbf{M}^{\top}\mathbf{M})}{\lambda_{\min^{+}}(\mathbf{M}^{\top}\mathbf{M})}}\right\rceil$. Define $\mathcal{P}_{\mathbf{M}}(t)=1-\frac{T_{n}\left((L_{\mathbf{M}}+\mu_{\mathbf{M}}-2t)/(L_{\mathbf{M}}-\mu_{\mathbf{M}})\right)}{T_{n}\left((L_{\mathbf{M}}+\mu_{\mathbf{M}})/(L_{\mathbf{M}}-\mu_{\mathbf{M}})\right)}$, where $T_{n}$ is the Chebyshev polynomial of the first kind of degree $n$ defined by $T_{n}(t)=\frac{1}{2}\left(\left(t+\sqrt{t^{2}-1}\right)^{n}+\left(t-\sqrt{t^{2}-1}\right)^{n}\right)$. Then, $\mathcal{P}_{\mathbf{M}}(0)=0$, and
@@
Results of this section are summarized in the following Lemma 4.
@@
$\mathbf{W^{\prime}}=P_{\mathbf{W}}(\mathbf{W})=\mathcal{P}_{\sqrt{\mathbf{W}}}(\mathbf{W})$ (19)
@@
$\mathbf{K}=\sqrt{P_{\mathbf{B}}(\mathbf{B}^{\top}\mathbf{B})}=\sqrt{\mathcal{P}_{\mathbf{B}}(\mathbf{B}^{\top}\mathbf{B})}.$ (20)
@@ 4.3 Base algorithm
Our base algorithm, Algorithm 1, is the Proximal Alternating Predictor-Corrector (PAPC) with Nesterov’s acceleration, called Accelerated PAPC (APAPC). It was proposed in Salim et al. (2022a) to obtain an optimal algorithm for optimization problems formulated as (21). See Kovalev et al. (2020); Salim et al. (2022b) for the review of related algorithms and history of their development.
APAPC algorithm formulates as Algorithm 1, and its convergence properties are given in Proposition 1.
@@
Assume that the matrix $\mathbf{K}$ in (21) satisfies $\mu_{\mathbf{K}}>0$ and $\mathbf{b}^{\prime}\in\operatorname{\mathrm{range}}\mathbf{K}$, and denote $\kappa_{\mathbf{K}}=\frac{L_{\mathbf{K}}}{\mu_{\mathbf{K}}}$. Also assume that the function $G$ is $L_{G}$-smooth and $\mu_{G}$-strongly convex. Set the parameter values of Algorithm 1 as $\tau=\min\left\{1,\frac{1}{2}\sqrt{\frac{\kappa_{\mathbf{K}}}{\kappa_{G}}}\right\}$, $\eta=\frac{1}{4\tau L_{G}}$, $\theta=\frac{1}{\eta L_{\mathbf{K}}}$ and $\alpha=\mu_{G}$. Denote by $u^{*}$ the solution of problem (21) and by $z^{*}$ the solution of its dual problem satisfying $z^{*}\in\operatorname{\mathrm{range}}\mathbf{K}$. Then the iterates $u^{k},z^{k}$ of Algorithm 1 satisfy

### v4

@@ 1 Introduction
Coupled constraints arise in various application scenarios, where sharing resources or information takes place. Often, due to the distributed nature of such problems, decentralization is desired for communication and/or privacy related reasons. Let us briefly describe several practical cases of optimization problems with coupled constraints.
@@
To handle the consensus constraint, decentralized algorithms either reformulate it as $\sum_{i=1}^{n}\mathbf{W}_{i}x_{i}=0$, where $\mathbf{W}_{i}$ is the $i$-th vertical block of a gossip matrix (an example of which is the communication graph’s Laplacian), or utilize the closely related mixing matrix approach Gorbunov et al. (2022). Mixing and gossip matrices are used because they are communication-friendly: calculating the sum $\sum_{i=1}^{n}\mathbf{W}_{i}x_{i}$ only requires each compute node to communicate once with each of its adjacent nodes. Clearly, consensus optimization with gossip matrix reformulation can be reduced to (1) by setting $\mathbf{A}_{i}=\mathbf{W}_{i}$. However, the principal difference between this example and (1), is that (1) does not assume $\mathbf{A}_{i}$ to be communication-friendly. We discuss the complexity of the reduction from consensus optimization to optimization with coupled constraints in Appendix˜A.
@@ 2 Related work and our contribution
Decentralized optimization algorithms were initially proposed for consensus optimization Nedić and Ozdaglar (2009), based on earlier research in distributed optimization Tsitsiklis (1984); Bertsekas and Tsitsiklis (1989) and algorithms for decentralized averaging (consensus or gossip algorithms) Boyd et al. (2006); Olshevsky and Tsitsiklis (2009), which assumed the existence of a communication network, as does the present paper. The optimal complexity for consensus optimization was first achieved with a dual accelerated gradient descent in Scaman et al. (2017), where the method required computing gradients of Fenchel conjugates of $f_{i}(x)$. The corresponding complexity lower bounds were also established in the same paper. This result was later generalized to primal algorithms (which use gradients of the functions $f_{i}(x)$ themselves) Kovalev et al. (2020), time-varying communication graphs Li and Lin (2021); Kovalev et al. (2021) and methods that use stochastic gradients Dvinskikh and Gasnikov (2021). Today there also exist algorithms with communication compression Beznosikov et al. (2023), asynchronous algorithms Koloskova (2024), algorithms for saddle-point formulations Rogozin et al. (2021) and gradient-free oracles Beznosikov et al. (2020), making decentralized consensus optimization a quite well-developed field Nedić (2020); Gorbunov et al. (2022), benefiting systems control Ram et al. (2009) and machine learning Lian et al. (2017).
Beginning with the addition of local constraints to consensus optimization Nedic et al. (2010); Zhu and Martinez (2011), constrained decentralized optimization has been established as a research direction. A zoo of distributed problems with constraints was investigated in Necoara et al. (2011); Necoara and Nedelcu (2014; 2015).
Primarily motivated by the demand from the power systems community, various decentralized algorithms for coupled constraints have been proposed. Generally designed for versatile engineering applications, many of these algorithms assume restricted function domains Wang and Hu (2022); Liang et al. (2019); Nedić et al. (2018); Gong and Zhang (2023); Zhang et al. (2021); Wu et al. (2022), nonlinear inequality constraints Liang et al. (2019); Gong and Zhang (2023); Wu et al. (2022), time-varying graphs Zhang et al. (2021); Nedić et al. (2018) or utilize specific problem structure Wang and Hu (2022).
@@
Works of Doan and Olshevsky (2017); Li et al. (2018); Nedić et al. (2018) focus on the resource allocation problem. For undirected time-varying graphs Doan and Olshevsky (2017) proposes a first-order algorithm with $O(B\kappa_{f}n^{2}\ln\frac{1}{\varepsilon})$ communication and derivative computation complexity bound, where $B$ is the time required for the time-varying graph to reach connectivity. Li et al. (2018) applies a combination of gradient tracking and push-sum approaches from Nedic et al. (2017) to obtain linear convergence on directed time-varying graphs in the restricted domain case, i.e., $x_{i}\in\Omega_{i}$, where $\Omega_{i}$ is a nonempty closed convex set. Nedić et al. (2018) achieves accelerated linear convergence via a proximal point method in the restricted domain case. When $\Omega_{i}=\mathbb{R}^{d}$, they also show that Nesterov’s accelerated gradient descent can be applied to achieve optimal $O(\sqrt{\kappa_{\mathbf{W}}}\sqrt{\kappa_{f}}\ln\frac{1}{\varepsilon})$ communication complexity. In Gong and Zhang (2023) an inexact proximal-point method is proposed to solve problems with coupled affine equality and convex inequality constraints. Linear convergence is proved when the inequalities are absent, and $\Omega_{i}$ are convex polyhedrons. The papers Wu et al. (2022), Chang (2016), Falsone et al. (2020) present algorithms with sub-linear convergence.
As summarized in Table 1, no accelerated linearly convergent algorithms for general affine-equality coupled constraints were present in the literature prior to our work. Also, most of the algorithms require proximal oracle, which allows to handle more general problem formulations, but has higher computational burden than the first-order oracle. We propose a new first-order decentralized algorithm with optimal (accelerated) linear convergence rate. We prove its optimality by providing lower bounds for the number of objective’s gradient computations, matrix multiplications and decentralized communications, which match complexity bounds for our algorithm.
@@ 3 Mathematical setting and assumptions
By $\kappa_{f}$ we denote the condition number $\kappa_{f}={L_{f}}/{\mu_{f}}$.
@@
where the matrix $\mathbf{S}\in\mathbb{R}^{m\times m}$ is defined as $\mathbf{S}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{A}_{i}\mathbf{A}_{i}^{\top}$. We also define the condition number of the block-diagonal matrix $\mathbf{A}=\mathrm{diag}\left(\mathbf{A}_{1},\ldots,\mathbf{A}_{n}\right)\in\mathbb{R}^{mn\times d}$ as $\kappa_{\mathbf{A}}=L_{\mathbf{A}}/\mu_{\mathbf{A}}$.
@@
We also assume the existence of a so-called gossip matrix $W\in\mathbb{R}^{n\times n}$ associated with the communication network $\mathcal{G}$, which satisfies the following assumption.
@@ 4.1 Strongly convex communication-friendly reformulation
and multiplication of a vector $y=(y_{1},\ldots,y_{n})\in(\mathbb{R}^{m})^{n}$ by $\mathbf{W^{\prime}}$ can be performed efficiently in the decentralized manner if its $i$-th block component $y_{i}$ is stored at $i$-th node of the computation network. Similarly to eq.˜6, we define
@@
Due to the definition of $\mathbf{W}$ and eq.˜7, the simplest choice for $\mathbf{W^{\prime}}$ might be to set $\mathbf{W^{\prime}}=\mathbf{W}$. Later we will specify another way to choose $\mathbf{W^{\prime}}$ for optimal algorithmic performance.
@@
From the definitions of $\mathbf{A}$, $\mathbf{b}$ and $\mathcal{L}_{m}^{\bot}$ (eq.˜3) it is clear that $\sum_{i=1}^{n}(\mathbf{A}_{i}x_{i}-b_{i})=0$ if and only if $\mathbf{A}x-\mathbf{b}\in\mathcal{L}_{m}^{\perp}$. Since $\operatorname{\mathrm{range}}\mathbf{W^{\prime}}=\mathcal{L}^{\perp}_{m}$, the constraint in problem (10) is equivalent to the coupled constraint in (1). For all $x,y$ satisfying the constraint, the augmented objective function $G(x,y)$ is equal to the original objective function $F(x)$. Therefore, problem (10) is equivalent to problem (1).
The following Lemma˜1 shows that the function $G(x,y)$ is strongly convex and smooth.
@@
Let the matrix $\mathbf{B}\in\mathbb{R}^{mn\times(d+mn)}$ be defined as $\mathbf{B}=\left[\begin{matrix}\mathbf{A}&\gamma\mathbf{W^{\prime}}\end{matrix}\right]$. The following Lemma˜2 connects the spectral properties of $\mathbf{B}$, $\mathbf{A}$ and $\mathbf{W^{\prime}}$.
@@
Proofs of Lemma˜1 and Lemma˜2 are provided in Appendix B.
@@ 4.2 Chebyshev acceleration
Chebyshev acceleration allows us to decouple the number of computations of the objective’s gradient $\nabla F(x)$ from the properties of the communication network and the constraint matrix — specifically, from the condition numbers $\kappa_{\mathbf{W}}$ and $\kappa_{\mathbf{A}}$. The Chebyshev trick enables to replace the matrix with a matrix polynomial with a better condition number.
Consider some affine relation $\mathbf{M}u=\mathbf{d}$ and let $\mathcal{P}_{\mathbf{M}}$ be a polynomial such that $\mathcal{P}_{\mathbf{M}}(\lambda)=0\Leftrightarrow\lambda=0$ for any eigenvalue $\lambda$ of $\mathbf{M}^{\top}\mathbf{M}$. Note that here we interchangeably use $\mathcal{P}$ as a polynomial of a matrix and a polynomial of a scalar. We denote any feasible point for the constraint $\mathbf{M}u=\mathbf{d}$ as $u_{0}$. Then,
$\mathbf{M}u=\mathbf{d}$ $\Leftrightarrow\mathbf{M}(u-u_{0})=0\overset{(a)}{\Leftrightarrow}\mathbf{M}^{\top}\mathbf{M}(u-u_{0})=0$
$\overset{(b)}{\Leftrightarrow}\mathcal{P}_{\mathbf{M}}(\mathbf{M}^{\top}\mathbf{M})(u-u_{0})=0\overset{(c)}{\Leftrightarrow}\sqrt{\mathcal{P}_{\mathbf{M}}(\mathbf{M}^{\top}\mathbf{M})}(u-u_{0})=0$
where (a) and (c) is due to $\ker\mathbf{M}^{\top}\mathbf{M}=\ker\mathbf{M}$; (b) is due to $\ker\mathcal{P}_{\mathbf{M}}(\mathbf{M}^{\top}\mathbf{M})=\ker\mathbf{M}^{\top}\mathbf{M}$ by the assumption about $\mathcal{P}_{\mathbf{M}}(\lambda)$.
Following Salim et al. (2022a) and Scaman et al. (2017), we use the translated and scaled Chebyshev polynomials, because they are the best at compressing the spectrum Auzinger and Melenk (2011).
@@
Consider a matrix $\mathbf{M}$. Let $\ell=\left\lceil\sqrt{\frac{L_{\mathbf{M}}}{\mu_{\mathbf{M}}}}\right\rceil\geq\left\lceil\sqrt{\frac{\lambda_{\max}(\mathbf{M}^{\top}\mathbf{M})}{\lambda_{\min^{+}}(\mathbf{M}^{\top}\mathbf{M})}}\right\rceil$. Define $\mathcal{P}_{\mathbf{M}}(t)=1-\frac{T_{\ell}\left((L_{\mathbf{M}}+\mu_{\mathbf{M}}-2t)/(L_{\mathbf{M}}-\mu_{\mathbf{M}})\right)}{T_{\ell}\left((L_{\mathbf{M}}+\mu_{\mathbf{M}})/(L_{\mathbf{M}}-\mu_{\mathbf{M}})\right)}$, where $T_{\ell}$ is the Chebyshev polynomial of the first kind of degree $n$ defined by $T_{\ell}(t)=\frac{1}{2}\left(\left(t+\sqrt{t^{2}-1}\right)^{\ell}+\left(t-\sqrt{t^{2}-1}\right)^{\ell}\right)$. Then, $\mathcal{P}_{\mathbf{M}}(0)=0$, and
@@
Results of this section are summarized in the following Lemma˜4.
@@
$\mathbf{W^{\prime}}=\mathcal{P}_{\sqrt{\mathbf{W}}}(\mathbf{W})$ (18)
@@
$\mathbf{K}=\sqrt{\mathcal{P}_{\mathbf{B}}(\mathbf{B}^{\top}\mathbf{B})}.$ (19)
@@ 4.3 Base algorithm
Our base algorithm, Algorithm˜1, is the Proximal Alternating Predictor-Corrector (PAPC) with Nesterov’s acceleration, called Accelerated PAPC (APAPC). It was proposed in Salim et al. (2022a) to obtain an optimal algorithm for optimization problems formulated as (20). See Kovalev et al. (2020); Salim et al. (2022b) for the review of related algorithms and history of their development.
APAPC algorithm formulates as Algorithm˜1, and its convergence properties are given in Proposition˜1.
@@
Assume that the matrix $\mathbf{K}$ in (20) satisfies $\mu_{\mathbf{K}}>0$ and $\mathbf{b}^{\prime}\in\operatorname{\mathrm{range}}\mathbf{K}$, and denote $\kappa_{\mathbf{K}}=\frac{L_{\mathbf{K}}}{\mu_{\mathbf{K}}}$. Also assume that the function $G$ is $L_{G}$-smooth and $\mu_{G}$-strongly convex. Set the parameter values of Algorithm˜1 as $\tau=\min\left\{1,\frac{1}{2}\sqrt{\frac{\kappa_{\mathbf{K}}}{\kappa_{G}}}\right\}$, $\eta=\frac{1}{4\tau L_{G}}$, $\theta=\frac{1}{\eta L_{\mathbf{K}}}$ and $\alpha=\mu_{G}$. Denote by $u^{*}$ the solution of problem (20) and by $z^{*}$ the solution of its dual problem satisfying $z^{*}\in\operatorname{\mathrm{range}}\mathbf{K}$. Then the iterates $u^{k},z^{k}$ of Algorithm˜1 satisfy